# Learning Safe Strategies for Value Maximizing Buyers in Uniform Price Auctions

### v2

@@ 5.  Competing Against Richer Classes of Bidding Strategies
In previous sections, we characterized the class of safe bidding strategies and proposed a learning algorithm that obtain sublinear regret, where the regret is computed against a clairvoyant that also selects the optimal safe bidding strategy. In this section, we consider the cases where the clairvoyant can choose the optimal strategy from richer bidding classes, which we will describe shortly. We show that under Algorithm 1, the class of safe bidding strategies achieves sublinear (approximate) regret when competing against these stronger benchmarks. The performance metric we consider in this section is
$\alpha\text{-}\mathsf{REG}_{\mathscr{B}_{c}}=\alpha\cdot\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\,.$
Here, $\mathscr{B}_{c}$ is the class of bidding strategies from which the clairvoyant (which serves as our benchmark) chooses the optimal bidding strategy whereas the bidder (learner) submits bids, $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}$, in each round unless stated otherwise.
@@
In prior works on approximate regret, $\alpha$ typically measures the hardness of the offline problem, such as $\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (see Streeter and Golovin (2008); Niazadeh et al. (2022)). In contrast, our work reinterprets $\alpha$ as richness ratio, capturing the relative richness of the clairvoyant’s strategy class compared to the learner’s. Specifically, it quantifies how closely the value achieved by the optimal strategy in $\mathscr{U}_{m}^{\star}$ approximates that by the optimal strategy in $\mathscr{B}_{c}$. If $\mathscr{B}_{c}=\mathscr{U}_{m}^{\star}$, then $\alpha=1$ (see Section 4). By considering a richer class of strategies for the clairvoyant, we aim to quantify the robustness and performance of safe strategies against stronger benchmarks.
@@
for $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$. In words, $\Lambda_{\mathscr{B}_{c}}$ is the maximum ratio of the value obtained by the optimal bidding strategy in $\mathscr{B}_{c}$ and that by the optimal strategy in $\mathscr{U}_{m}^{\star}$ over all valuation curves and bid histories in an offline setting. We drop the argument $\mathbf{v}$ from $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ when the context is clear.
@@
Suppose the clairvoyant chooses the optimal strategy from the class $\mathscr{B}_{c}$ such that $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}\leq\lambda$ and this bound is tight. Then, Algorithm 1 obtains $\frac{1}{\lambda}$-$\mathsf{REG}_{\mathscr{B}_{c}}\leq O(M\sqrt{mT\log M})$ under full information feedback and $\frac{1}{\lambda}$-$\mathsf{REG}_{\mathscr{B}_{c}}\leq O(M\sqrt{m^{3}MT\log M})$ under bandit feedback.
@@
$\frac{1}{\lambda}\text{-}\mathsf{REG}_{\mathscr{B}_{c}}$ $=\frac{1}{\lambda}\cdot\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]$
$\leq\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]=\mathsf{REG}\,.\vspace{-0.5cm}$
By Theorem 4.3, we get the stated regret upper bound. ∎
5.1.  Richness Ratio of $m$-uniform Non-Safe Strategies
Here, we consider the clairvoyant class $\mathscr{B}_{c}$ to be the class of bidding strategies containing at most $m$ bid-quantity pairs, which are feasible only for the given bid history $\mathcal{H}_{-}$, rather than for every possible sequence of competing bids as in the case of safe strategies. The following theorem characterizes the richness ratio of this class, denoted as $\mathscr{F}^{\mathcal{H}_{-}}_{m}$:
@@ 5.1.  Richness Ratio of $m$-uniform Non-Safe Strategies
Theorem 5.3 implies that restricting the learner to safe bidding strategies has a bounded cost and does not lead to an arbitrary loss in the obtained value as the upper bound is independent of $m$. Building on this, Corollary 5.2 shows that Algorithm 1 results in $\frac{1}{2}$-approximate sublinear regret in the online setting when the safe strategies compete against $\mathscr{F}^{\mathcal{H}_{-}}_{m}$. Additionally, the factor-of-two loss represents a worst-case scenario, occurring in a highly non-trivial setting (see Section D.2). In practice, we expect these strategies to perform near-optimally, a claim further supported by the numerical simulations presented in Section 5.3.
Proof Sketch of Theorem 5.3 (Upper Bound). We now present the central ideas to prove the upper bound, which are also used to show the subsequent theorems, albeit with some changes specific to each case. Fixing a bid history, $\mathcal{H}_{-}$, we consider the expression in Section 5. Maximizing over all bid histories and valuation curves, we get the desired result.
Firstly, we bridge the we bridge the gap between the bidding strategies of the clairvoyant and the learner by constructing a restricted class of safe strategies, $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})\subseteq\mathscr{U}_{m}^{\star}$. This restricted class is exponentially smaller in size and is derived based on the optimal strategy selected by the clairvoyant (see exact characterization in Eq. 17 for Theorem 5.3 and Algorithm 3 for Theorem 5.4). We demonstrate that considering this restricted class of safe strategies is sufficient to derive tight upper bounds.
Second, as a crucial step, we show that for any $m$-uniform safe strategy $\mathbf{b}=\langle(w_{Q_{1}},q_{1}),\dots,(w_{Q_{m}},q_{m})\rangle$,
$V(\mathbf{b};\bm{\beta}_{-})=\max_{\ell\in[m]}V((w_{Q_{\ell}},Q_{\ell});\bm{\beta}_{-})$
for all competing bids $\bm{\beta}_{-}$, where, $Q_{\ell}=\sum_{j=1}^{\ell}q_{j}$. Thus, the value obtained by any $m$-uniform safe strategy can be determined by knowing the value obtained by the strategies in $\mathscr{U}_{1}^{\star}$, allowing us to focus solely on 1-uniform strategies.
As the third key idea, we partition the $T$ rounds into sets $T_{j}$, based on the least winning bid of the optimal bidding strategy chosen by the clairvoyant (see details in Eq. 15) and derive lower bounds on the value obtained by 1-uniform strategies in $\mathscr{U}_{m}^{\star}(\mathcal{H}_{-})$ in terms of the value obtained by optimal strategy chosen by the clairvoyant in each partition $T_{j}$. With these three key ideas, we bound the ratio in Section 5. The detailed proof is presented in Section C.1.
Proof Sketch of Theorem 5.3 (Tight Lower Bound). For any $\delta\in(0,1]$, we construct a tight lower-bound instance for any $m\in\mathbb{N}$ by considering an offline problem where $M=T=N^{2m}$ and $K=M+1$, with $N=O(1/\delta)$. We set $\mathbf{v}=[1,v,\dots,v]\in\mathbb{R}^{M}$, where $v=1-O(\delta)$. The sequence of $T$ auctions is divided into $2m$ partitions of exponentially varying sizes, each with an identical competing bid profile. In each bid profile, there are two types of bids: ‘big’ and ‘small’. All ‘big’ bids are set to $C\gg 1$, while the ‘small’ bids remain identical within a partition. The partition sizes, the number of ‘small’ competing bids, and the exact values of these bids are carefully chosen to ensure that $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-};\mathbf{v})\geq 2-\delta$ (see Sections D.1 to D.2).
5.2.  Richness Ratio of $m^{\prime}$-uniform Strategies
We first consider the case when the clairvoyant selects the optimal strategy from the class of safe bidding strategies with at most $m^{\prime}$ bid-quantity pairs where $m^{\prime}\geq m$.
@@ 5.2.  Richness Ratio of $m^{\prime}$-uniform Strategies
For any $m,m^{\prime}\in\mathbb{N}$ such that $m^{\prime}\geq m$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}\leq\frac{m^{\prime}}{m}$. Additionally, there exist $\mathcal{H}_{-}$ and $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{m^{\prime}}{m}-\delta$. Thus, the richness ratio of the class $\mathscr{U}_{m^{\prime}}^{\star}$ is $\alpha=m/m^{\prime}$.
Suppose the clairvoyant can now select the optimal strategy from $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$, the class of bidding strategies with at most $m^{\prime}$ bid-quantity pairs ($m^{\prime}\geq m$) that are feasible for the bid history $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$, but not necessarily safe. Then,
@@
For any $m,m^{\prime}\in\mathbb{N}$ such that $m^{\prime}\geq m$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}\leq\frac{2m^{\prime}}{m}$. Moreover, there exist $\mathcal{H}_{-}$ and $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{2m^{\prime}}{m}-\delta$. Thus, the richness ratio of the class $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$ is $\alpha=m/2m^{\prime}$.
@@
By Corollary 5.2, for $m^{\prime}\geq m$, Algorithm 1 obtains $\frac{m}{m^{\prime}}$-approximate (resp. $\frac{m}{2m^{\prime}}$-approximate) sublinear regret in the online setting when the clairvoyant draws the optimal strategy from the class of safe (resp. feasible but not necessarily safe) strategies with at most $m^{\prime}$ bid-quantity pairs.
@@
The richness ratio in Theorem 5.4 and Theorem 5.5 can theoretically become arbitrarily small as $m^{\prime}$ grows, for a fixed $m$. However, the bidder has the flexibility to choose $m$; increasing $m$, though, incurs higher space and time complexity for Algorithm 1. Moreover, for any given pair $(m,m^{\prime})$, the bound is tight under a highly non-trivial setting (see details in Sections D.3 to D.4). In practice, as shown in Section 5.3, the observed richness ratios are substantially better than the theoretical bounds, suggesting that the worst-case scenarios are unlikely to arise in real-world applications.
5.3.  Estimating The Richness Ratio: A Case Study
In this section, we estimate the richness ratio (equivalently $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}$) for the discussed classes of bidding strategies by experiments using EU ETS emission permit auction data from 2022 and 2023. Detailed methods are provided in Appendix E, and the results are shown in Fig. 4.
@@ 5.3.  Estimating The Richness Ratio: A Case Study
Figure 4: The left (resp. right) figure refers to (the upper bound on) $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ (resp. $\Lambda_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-})$) as a function of $m$.
@@
Estimating $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$. The left plot of Fig. 4 shows that (the upper bound for) $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})$ is significantly better than the theoretical bound of 2. The decaying trend of the plot with increasing $m$ can be attributed to the fact that we are computing an upper bound for the value achieved by the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m}$, which is independent of $m$, rather than the exact value (see Section E.2). For $m\geq 4$, even the upper bound of $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-})\sim 1.05$ indicating that the safe bidding strategies are near-optimal.

### v4

@@ 5. Competing Against Richer Classes of Bidding Strategies
In previous sections, we characterized the class of safe bidding strategies and proposed a learning algorithm that obtains sublinear regret, where the regret is computed against a clairvoyant that also selects the optimal safe bidding strategy. In this section, we consider the cases where the clairvoyant can choose the optimal strategy from richer bidding classes, which we will describe shortly. We show that the bidder who follows safe bidding strategies and runs Algorithm 1 in the full information setting achieves sublinear (approximate) regret when competing against these stronger benchmarks. A similar result also holds in the bandit setting, mutatis mutandis. The performance metric we consider in this section is
$\alpha\text{-}\textsf{REG}_{\mathscr{B}_{c}}=\alpha\cdot\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]\,.$ (8)
Here, $\mathscr{B}_{c}$ is the class of bidding strategies from which the clairvoyant (which serves as our benchmark) chooses the optimal bidding strategy whereas the bidder (learner) submits bids, $\mathbf{b}^{t}\in\mathscr{U}_{m}^{\star}$, in each round unless stated otherwise. Here, $\alpha\in(0,1]$ is defined as the richness ratio.
@@
In prior works on approximate regret, the parameter $\alpha$ typically measures the hardness of the offline problem, such as $\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})$ (see Streeter and Golovin (2008); Niazadeh et al. (2022)). In contrast, our work reinterprets $\alpha$ as richness ratio, capturing the relative richness of the clairvoyant’s strategy class compared to the learner’s. Specifically, it quantifies how closely the value achieved by the optimal safe strategy in $\mathscr{U}_{m}^{\star}$ approximates that by the optimal strategy in $\mathscr{B}_{c}$. If $\mathscr{B}_{c}=\mathscr{U}_{m}^{\star}$, then $\alpha=1$ (see Section 4). By considering a richer class of strategies for the clairvoyant, we aim to quantify the robustness and performance of safe strategies against stronger benchmarks.
@@
for $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$. In words, $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}$ is the maximum ratio of the value obtained by the optimal bidding strategy in $\mathscr{B}_{c}$ and that by the optimal strategy in $\mathscr{U}_{m}^{\star}$ over all valuation curves and all bid histories of arbitrary length.
@@
Suppose the clairvoyant chooses the optimal strategy from the class $\mathscr{B}_{c}$ such that $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}\leq\lambda$ and this bound is tight implying the richness ratio of $\mathscr{B}_{c}$ is $\alpha=\frac{1}{\lambda}$. Then, Algorithm 1 obtains $\alpha$-$\textsf{REG}_{\mathscr{B}_{c}}\leq O(M\sqrt{mT\log M})$ in the full information setting.
@@
$\alpha\text{-}\textsf{REG}_{\mathscr{B}_{c}}$ $=\frac{1}{\lambda}\cdot\max_{\mathbf{b}\in\mathscr{B}_{c}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]$
$\leq\max_{\mathbf{b}\in\mathscr{U}_{m}^{\star}}\sum_{t=1}^{T}V(\mathbf{b};\bm{\beta}_{-}^{t})-\sum_{t=1}^{T}\mathbb{E}[V(\mathbf{b}^{t};\bm{\beta}_{-}^{t})]=\textsf{REG},$
where the first equality holds as $\alpha=\frac{1}{\lambda}$. By Theorem 4.3, we get the stated regret upper bound. ∎
Recall that we defined richness of a class of bidding strategies along two dimensions: (i) the safety of strategies (safe vs. only RoI-feasible), and (ii) expressiveness, i.e., the number of bid-quantity pairs allowed. In the subsequent sections, we consider classes of strategies that are richer along either one or both of the said dimensions and compute the corresponding richness ratios, $\alpha$.
5.1. The Class of RoI-feasible Strategies with $\leq m$ Bid-Quantity Pairs
First, we consider the clairvoyant class $\mathscr{B}_{c}$ to be the class of strategies containing at most $m$ bid-quantity pairs, which are RoI-feasible only for the given bid history $\mathcal{H}_{-}$, rather than for every possible sequence of competing bids as in the case of safe strategies. We denote this class by $\mathscr{F}^{\mathcal{H}_{-}}_{m}$. Then,
@@ 5.1. The Class of RoI-feasible Strategies with $\leq m$ Bid-Quantity Pairs
We present the proof of Theorem 5.3 in Section 5.4. Theorem 5.3 implies that restricting the learner to safe bidding strategies has a bounded cost and does not lead to an arbitrary loss in the obtained value as the upper bound is independent of $m$. Building on this, Corollary 5.2 shows that Algorithm 1 results in $\frac{1}{2}$-approximate sublinear regret in the online setting when the safe strategies compete against $\mathscr{F}^{\mathcal{H}_{-}}_{m}$. Additionally, the factor-of-two loss represents a worst-case scenario, occurring in a highly non-trivial setting (see Section D.1). In practice, we expect these strategies to perform near-optimally, a claim further supported by the numerical simulations presented in Section 5.3.
5.2. Classes of Strategies with $\leq m^{\prime}$ Bid-Quantity Pairs For $m^{\prime}\geq m$
We now consider the setting where the clairvoyant selects the optimal strategy from the class of safe strategies with at most $m$’ bid-quantity pairs, where $m’\geq m$. This strategy class, $\mathscr{U}_{m’}^{\star}$, is richer than the bidder’s class $\mathscr{U}_{m}^{\star}$, although, along the expressiveness dimension compared to the class $\mathscr{F}^{\mathcal{H}_{-}}_{m}$ discussed in the previous section, which is richer along the safety dimension. Then,
@@ 5.2. Classes of Strategies with $\leq m^{\prime}$ Bid-Quantity Pairs For $m^{\prime}\geq m$
For any $m,m^{\prime}\in\mathbb{N}$ such that $m^{\prime}\geq m$, $\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}\leq\frac{m^{\prime}}{m}$. Additionally, there exist $\mathcal{H}_{-}$ and $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$,
$\Lambda_{\mathscr{U}_{m^{\prime}}^{\star},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{m^{\prime}}{m}-\delta\,.$
Thus, the richness ratio of the class $\mathscr{U}_{m^{\prime}}^{\star}$ is $\alpha=m/m^{\prime}$.
Finally, suppose the clairvoyant can now select the optimal strategy from $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$, the class of bidding strategies with at most $m^{\prime}$ bid-quantity pairs ($m^{\prime}\geq m$) that are only RoI-feasible for the bid history $\mathcal{H}_{-}=[\bm{\beta}_{-}^{t}]_{t\in[T]}$. This strategy class is richer than the bidder’s strategy class along both dimensions. Then,
@@
For any $m,m^{\prime}\in\mathbb{N}$ such that $m^{\prime}\geq m$, $\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}\leq\frac{2m^{\prime}}{m}$. Moreover, there exist $\mathcal{H}_{-}$ and $\mathbf{v}$ such that for any $\delta\in(0,\frac{1}{2}]$, we have
$\Lambda_{\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})\geq\frac{2m^{\prime}}{m}-\delta\,.$
Thus, the richness ratio of the class $\mathscr{F}^{\mathcal{H}_{-}}_{m^{\prime}}$ is $\alpha=m/2m^{\prime}$.
@@
By Corollary 5.2, for $m^{\prime}\geq m$, Algorithm 1 obtains $\frac{m}{m^{\prime}}$-approximate (resp. $\frac{m}{2m^{\prime}}$-approximate) sublinear regret (same as Theorem 4.3) in the online setting when the clairvoyant draws the optimal strategy from the class of safe (resp. RoI-feasible) strategies with at most $m^{\prime}$ bid-quantity pairs.
@@
The richness ratio in Theorem 5.4 and Theorem 5.5 can theoretically become arbitrarily small as $m^{\prime}$ increases, for a fixed $m$. However, the bidder retains the flexibility to choose $m$: increasing $m$, though, incurs higher space and time complexity for Algorithm 1. Moreover, for any given pair $(m,m^{\prime})$, the bounds are tight for carefully constructed non-trivial hard instances (see details in Sections D.2 to D.3). In practice, as shown in Section 5.3, the empirical richness ratios are substantially better than the worst-case bounds, suggesting that such pathological scenarios are unlikely to arise in real-world settings.
We conclude the section with two key observations. First, when $m=m$’, Theorem 5.4 implies $\alpha=1$, which corresponds to the setting where both the bidder and the clairvoyant follow the class of safe strategies with at most $m$ bid-quantity pairs, the case considered in Section 4. Similarly, when $m=m$’, Theorem 5.5 yields $\alpha=\frac{1}{2}$, recovering the result of Theorem 5.3. Second, recall that we defined the richness of a strategy class along two dimensions: safety and expressiveness. The results in this section imply that when a benchmark strategy class is richer along both dimensions (Theorem 5.5), its overall richness ratio is the product of the richness ratios along each individual dimension (Theorem 5.3 and Theorem 5.4).
5.3. Computing Empirical Richness Ratio: A Case Study
As discussed in the previous section, constructing worst-case instances for showing tight bounds are significantly intricate that may not reflect real world settings. This motivates us to empirically estimate how well do the class of safe strategies approximate the richer strategy classes in practice. To this end, with slight abuse of notation, we define the empirical richness ratio as
$\alpha_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})=\frac{1}{\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})},$
where $\Lambda_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ is defined in Section 5. We assess the tightness of the theoretical bounds in practice by computing $\alpha_{\mathscr{B}_{c},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ using the EU ETS emission permit auction data from 2022 and 2023. However, only aggregate statistics of the submitted bids is publicly available for privacy reasons. Hence, we synthesize individual level bid data from these available statistics. The exact procedure to reconstruct the bid data is presented in Section F.1.
@@ 5.3. Computing Empirical Richness Ratio: A Case Study
Figure 4: The left (resp. right) figure refers to (the lower bound on) $\alpha_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ (resp. $\alpha_{\mathscr{U}_{m}^{\star},\mathscr{U}_{1}^{\star}}(\mathcal{H}_{-},\mathbf{v})$) as a function of $m$. The shaded area corresponds to one standard deviation.
@@
Estimating $\alpha_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$. Since computing the exact value attained by the optimal strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m}$ is non-trivial, we instead compute a uniform upper bound that is independent of $m$ (see Section F.2). The left plot of Fig. 4 shows that the empirical lower bound of $\alpha_{\mathscr{F}^{\mathcal{H}_{-}}_{m},\mathscr{U}_{m}^{\star}}(\mathcal{H}_{-},\mathbf{v})$ is substantially better than the theoretical worst-case bound of $1/2$. In particular, for $m\geq 4$, even the lower bound exceeds $0.95$, indicating that safe bidding strategies are nearly optimal in practice.